1.The length of the seconds hands of a clock is 1cm.The change in velocity in its tip is what?


2.A body starts moving with velocity u=10m/s.It experiences a retardation equal to 0.2v2 .Its velocity after 2s is what?

Grade:Upto college level

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Chandan das

given retardation = 0.2v2

or  -dv/dt =0.2v2

     -dv/v2=0.2 dt

 intigrate

1/V = 0.2t +c

now use initial condition  at t= o ,v=10 m/sec

so    c= 0.1

equation become 

 1/V = 0.2t +0.1


so at t=2 sec

1/v = 0.4+0.1

v= 2 m/sec

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin